Suppose that the lifetimes of electric light bulbs follow an Exponential distribution with mean

1000 hours. A

random sample of 40 light bulbs is tested. The probability that the mean

lifetime is at least 900 hours is:


1
Expert's answer
2022-06-06T16:36:51-0400

The Variance will be "Var(\bar{X})=\dfrac{\lambda^2}{n}."

"\mu=1000\ h, n=40"


"\lambda=\dfrac{1}{\mu}"

"\beta^2=Var(nX)"


"\beta=\sqrt{Var(X)}=\sqrt{\dfrac{\lambda^2}{n}}"

"=\sqrt{\dfrac{1}{40(1000)^2}}=1.58\times10^{-4}"

Exponential distribution


"P(X\ge900)=P(X>900)"

"=e^{-\beta(900)}=e^{-1.58\times10^{-4}(900)}=0.867448"